# Steering LVLMs via Sparse Autoencoder for Hallucination Mitigation

### v1

@@ Abstract
Large vision-language models (LVLMs) have achieved remarkable performance on multimodal tasks such as visual question answering (VQA) and image captioning. However, they still suffer from hallucinations, generating text inconsistent with visual input, posing significant risks in real-world applications. Existing approaches to address this issue focus on incorporating external knowledge bases, alignment training, or decoding strategies, all of which require substantial computational cost and time. Recent works try to explore more efficient alternatives by adjusting LVLMs’ internal representations. Although promising, these methods may cause hallucinations to be insufficiently suppressed or lead to excessive interventions that negatively affect normal semantics. In this work, we leverage sparse autoencoders (SAEs) to identify semantic directions closely associated with either hallucinations or actuality, realizing more precise and direct hallucination-related representations. Our analysis demonstrates that interventions along the faithful direction we identified can mitigate hallucinations, while those along the hallucinatory direction can exacerbate them. Building on these insights, we propose Steering LVLMs via SAE Latent Directions (SSL), a training-free method based on SAE-derived latent directions to mitigate hallucinations in LVLMs. Extensive experiments demonstrate that SSL significantly outperforms existing decoding approaches in mitigating hallucinations, while maintaining transferability across different model architectures with negligible additional time overhead.
@@
Zhenglin Hua1, Jinghan He2,3, Zijun Yao4, Tianxu Han5, Haiyun Guo2,3, Yuheng Jia1, Junfeng Fang6 1School of Computer Science and Engineering, Southeast University 2Foundation Model Research Center, Institute of Automation, Chinese Academy of Sciences 3School of Artificial Intelligence, University of Chinese Academy of Sciences 4Department of Computer Science and Technology, Tsinghua University 5Wuhan University of Technology 6National University of Singapore huazhenglin2003@gmail.com, hejinghan2022@ia.ac.cn
@@ 1 Introduction
We note the success of SAEs in extracting fine-grained semantic representations—specifically capturing whether the model knows certain entities—of abstract concepts in the field of large language models (LLMs) Ferrando et al. (2025). Inspired by this work, we extend the application of SAE-based analysis from LLMs to LVLMs, aiming to understand and mitigate hallucinations more precisely and directly by identifying and steering the internal semantic directions associated with hallucinatory and faithful content. Notably, SAEs builds on the Linear Representation Hypothesis Park et al. (2023), which posits that internal model representations can be expressed as sparse combinations of interpretable semantic directions Tigges et al. (2024); Li et al. (2023a). Specifically, we leverage the SAE provided by Zhang et al. (2024a) to identify latent directions that closely correlate with hallucinatory semantics as well as those aligned with faithful content. As illustrated in Figure 1, targeted interventions along faithful directions suppress hallucinatory generations, while perturbations along hallucinatory directions tend to elicit more factually incorrect content.
Building on this insight, we propose Steering LVLMs via SAE Latent Directions (SSL), a training-free approach based on SAE-derived latent directions to mitigate hallucinations in LVLMs. During the visual feature merging stage, we inject faithful semantic directions to amplify grounded semantic features and improve image–text consistency. In the subsequent language generation stage, we reduce projection onto hallucinatory semantic directions, thereby reducing the risk of generating factually incorrect content. Remarkably, although the SAE was trained on the LLaVA-Next-8b model, the extracted hallucination and factuality directions generalize seamlessly to other architectures (e.g., LLaVA1.5-7b model Liu et al. (2024c) and InstructBLIP-7b model Dai et al. (2023a)). Experimental evaluation on established LVLM hallucination benchmarks shows that SSL outperforms existing decoding approaches, confirming its effectiveness and efficiency in hallucination reduction.
@@
  1. We leverage SAEs to identify semantic directions that are highly correlated with hallucinatory and faithful object words in the representation space of LVLMs.
  2. We propose SSL, a training-free method that injects factuality semantic directions during visual feature fusion to reinforce grounded content and suppresses hallucination directions during language generation to proactively mitigate hallucinatory outputs.
@@ 3 Method
In this work, we introduce SSL, a training-free method for steering LVLMs. Our method consists of two principal components: semantic directions identification and steering LVLMs via SAE latents.
@@ 3.1 Semantic Directions Identification
Finally, to quantitatively evaluate the predictive power of SAE-derived directions in distinguishing hallucinatory from faithful samples, we design a set of classification experiments based on logistic regression. The model takes the latent activation values—either individually or in combination—as input features for a binary hallucination classification task. As illustrated in Figure 4, the results demonstrate that the latent semantic directions extracted by the SAE are discriminative (see Appendix B for more experimental details). Furthermore, combining the hall latent and faithful latent as input features yields further performance improvements.
@@ 3.2 Steering LVLMs via SAE Latents
Input: Scaling factor $\gamma$; steering layer $l_{s}$; semantic directions $d_{\text{hall}}$, $d_{\text{faithful}}$; residual stream at layer $l_{s}$: $[X_{l_{s},s},X_{l_{s},t},X_{l_{s},v},X_{l_{s},o}^{<t}]$
@@ 4.1 LVLMs
$\text{CHAIR}_{S}$downarrow $\text{CHAIR}_{I}$downarrow Avg.Len $\text{CHAIR}_{S}$downarrow $\text{CHAIR}_{I}$downarrow Avg.Len $\text{CHAIR}_{S}$downarrow $\text{CHAIR}_{I}$downarrow Avg.Len
@@ 4.2 Benchmarks
$\text{CHAIR}_{S}$ =|$=\frac{\left|\{\text{captions w/ hallucinatory objects}\}\right|}{\left|\{\text{total captions}\}\right|},$ (9)
$\text{CHAIR}_{I}$ =|$=\frac{\left|\{\text{hallucinatory objects}\}\right|}{\left|\{\text{total mentioned objects}\}\right|}.$
@@ 4.4 Implementation Details
We set $\gamma$ to $0.6$, $0.8$, and $0.2$ for LLaVA-NeXT, LLaVA-1.5, and InstructBLIP, respectively, to balance effective mitigation of hallucination with minimizing the invasiveness of state interventions. SSL is applied at $16^{\text{th}}$ layer for LLaVA-NeXT, $31^{\text{th}}$ layer for LLaVA-1.5 and $8^{\text{th}}$ layer for InstructBLIP. We faithfully replicate all baseline methods, implementing them based on their open-source codebases and configuring them according to the hyperparameters reported in the original papers. All experimental results are obtained under consistent base model, prompt, and generation parameter settings to ensure a fair comparison. For all methods involving beam search, we set the max_new_token to $512$ and the beam size to $5$.
@@ 5.1 Ablation Study
We conduct an ablation study on LLaVA-NeXT-8b to examine the impact of applying guidance at different layers. Figure 5 shows that the choice of guidance layer significantly affects model performance, setting $\gamma$ to $0.8$. For LLaVA-NeXT-8b, applying SSL at middle layers yields more effective mitigation of hallucinations, with layer $15$ achieving the best performance. For further analysis on steering layer selection and scaling factor $\gamma$ settings across different models, please refer to Appendix C and D.
@@ 5.2 Further Analysis
$\text{CHAIR}_{S}$downarrow $\text{CHAIR}_{I}$downarrow
LLaVA-NeXT w/ ASP 25.20 6.46
@@
InstructBLIP w/ ASP 38.52 11.45
InstructBLIP w/ fixed $\alpha$ 45.40 12.70
@@
To further validate the effectiveness of the semantic directions we identified, we compare the CHAIR benchmark across three model architectures under three settings: original model state, SSL, and Reverse-SSL (see Appendix E for details on Reverse-SSL). As shown in Figure 6, across all three model architectures, applying Reverse-SSL significantly increases hallucinations, while standard SSL guidance consistently reduces hallucinations. To further clearly illustrate the effectiveness of SSL, we present several qualitative examples in Appendix F.
@@ 7 Conclusion
This work explores the relationship between the hallucination in LVLMs and their internal latent representations. We construct a residual stream dataset for hallucinatory and faithful object tokens, and use SAE to extract the semantic directions corresponding to hallucination and Factuality. Based on this insight, we propose SSL, a training-free method that amplifies true semantics while suppressing potential hallucinations. Extensive experiments demonstrate that SSL outperforms existing methods. Furthermore, although the SAE was trained on LLaVA-Next, the semantic directions it extracted generalize well across different model architectures, further showcasing the potential of SAE in understanding and controlling the internal semantic representations of models.
@@
Currently, the only fully open-source multi-modal SAE is provided by LLM-Labs, trained on the $25^{\text{th}}$ layer of the LLaVA-Next 8b model. As a result, our study does not include a comparison of SAEs trained on other model architectures across different multi-modal models. Future work can focus on training multi-modal SAEs on various architectures to investigate whether the findings from this study generalize across different models.
@@ Appendix B A Set of Classification Experiments Based on Logistic Regression
Hall latent: The dimension $hall$ with the highest correlation to hallucinated object words is selected, forming a one-dimensional feature:
@@
Faithful latent: The dimension $faithful$ most correlated with faithful object words is selected, forming:
@@
$\text{CHAIR}_{S}$downarrow $\text{CHAIR}_{I}$downarrow Avg.Len
@@
Appendix C Choice of Steering Layer
@@
Figures 1011 and 12 present the results of ablation studies investigating the effect of introducing SSL at individual layers of the LLaVA-Next ($\gamma=0.6$), LLaVA-1.5 ($\gamma=0.8)$ and InstructBLIP ($\gamma=0.2$) models, respectively. For LLaVA-Next, we observe that applying SSL at the middle layers more effectively mitigates hallucinations, consistent with the results shown in Figure 5. For LLaVA-1.5, we observe that applying SSL at either the layer1 or deeper layers consistently mitigates hallucination. This observation aligns closely with findings reported by Zhang et al. (2024b), He et al. (2024) and Chen et al. (2025), who also found that layer1 or deeper layers interventions can significantly reduce hallucination in LLaVA-1.5. In contrast, for InstructBLIP, introducing SSL at shallow layers yields more substantial improvements, while deeper layer interventions contribute less noticeably to performance. We hypothesize that this is attributable to architectural and training differences in InstructBLIP, specifically, its shallow layers may already perform substantial cross-modal alignment early in the pipeline, making early-stage semantic guidance more impactful on overall generation quality. A deeper analysis of the layer-specific mechanisms in different multimodal architectures is left for future work.
Input: Scaling factor $\gamma$; steering layer $l_{s}$; semantic directions $d_{\text{hall}}$, $d_{\text{faithful}}$; residual stream at layer $l_{s}$: $[X_{l_{s},s},X_{l_{s},t},X_{l_{s},v},X_{l_{s},o}^{<t}]$
@@
Appendix D Choice of Scaling Factor
Tables 56, and7 report the ablation results on the effect of the scaling factor $\gamma$ in the SSL. For the LLaVA series of models, setting $\gamma$ to 0.6 or 0.8 effectively reduces hallucinations, indicating that moderate levels of semantic intervention are beneficial. However, when $\gamma>1.0$, the supervision becomes overly aggressive, disrupting the model’s behavior and leading to abnormal hallucination metrics. In contrast, for InstructBLIP, the best performance is observed at $\gamma=0.2$, while larger values similarly result in performance degradation. These findings highlight the importance of carefully calibrating the intensity of semantic guidance to balance model control and generation quality across different architectures.
$\text{CHAIR}_{S}$downarrow $\text{CHAIR}_{I}$downarrow Avg.Len

### v2

@@ Abstract
Large vision-language models (LVLMs) have achieved remarkable performance on multimodal tasks. However, they still suffer from hallucinations, generating text inconsistent with visual input, posing significant risks in real-world applications. Existing approaches to address this issue focus on incorporating external knowledge bases, alignment training, or decoding strategies, all of which require substantial computational cost and time. Recent works try to explore more efficient alternatives by adjusting LVLMs’ internal representations. Although promising, these methods may cause hallucinations to be insufficiently suppressed or lead to excessive interventions that negatively affect normal semantics. In this work, we leverage sparse autoencoders (SAEs) to identify semantic directions closely associated with faithfulness or hallucination, extracting more precise and disentangled hallucination-related representations. Our analysis demonstrates that interventions along the identified faithful direction can mitigate hallucinations, while those along the hallucinatory direction can exacerbate them. Building on these insights, we propose Steering LVLMs via SAE Latent Directions (SSL), a plug-and-play method based on SAE-derived latent directions to mitigate hallucinations in LVLMs. Extensive experiments demonstrate that SSL significantly outperforms existing decoding approaches in mitigating hallucinations, while maintaining transferability across different model architectures with negligible additional time overhead. The code is available at https://github.com/huazhenglin2003/SSL.
@@
Zhenglin Hua1,2, Jinghan He3,4, Zijun Yao5, Tianxu Han6, Haiyun Guo3,4, Yuheng Jia1,2, Junfeng Fang7 1School of Computer Science and Engineering, Southeast University 2Key Laboratory of New Generation Artificial Intelligence Technology and Its Interdisciplinary Applications (Southeast University) 3Foundation Model Research Center, Institute of Automation, Chinese Academy of Sciences 4School of Artificial Intelligence, University of Chinese Academy of Sciences 5Department of Computer Science and Technology, Tsinghua University 6Wuhan University of Technology 7National University of Singapore huazhenglin2003@gmail.com, hejinghan2022@ia.ac.cn
Equal contribution. Corresponding author.
@@ 1 Introduction
Recently, SAEs have shown success in extracting fine-grained semantic representations—specifically capturing whether the model knows certain entities—of abstract concepts in the field of large language models (LLMs) Ferrando et al. (2025). SAEs builds on the Linear Representation Hypothesis Park et al. (2023), which posits that internal model representations can be expressed as sparse combinations of interpretable semantic directions Tigges et al. (2024); Li et al. (2023a). Building on this foundation, Zhang et al. (2024a) utilize SAEs to interpret LVLM representations and analyze open-set semantics. However, their work does not address hallucination modeling or intervention. Inspired by these works, we extend the application of SAE-based analysis from LLMs to LVLMs, with a particular focus on mitigating hallucinations. Specifically, we leverage the SAE provided by Zhang et al. (2024a) to identify latent directions correlated with hallucinatory semantics as well as those aligned with faithful content, and steer the internal semantic representations of LVLMs, aiming to understand and mitigate hallucinations more precisely and directly. As illustrated in Figure 1, targeted interventions along faithful directions suppress hallucinatory generations, while perturbations along hallucinatory directions tend to elicit more factually incorrect content.
Building on this insight, we propose Steering LVLMs via SAE Latent Directions (SSL), a plug-and-play approach based on SAE-derived latent directions to mitigate hallucinations in LVLMs. During the visual feature merging stage, we inject faithful semantic directions to amplify grounded semantic features and improve image–text consistency. In the subsequent language generation stage, we reduce projection onto hallucinatory semantic directions, thereby reducing the risk of generating factually incorrect content. Remarkably, although the SAE was trained on the LLaVA-Next-8b model, the extracted hallucination and factuality directions generalize seamlessly to other architectures (e.g., LLaVA1.5-7b model Liu et al. (2024c) and InstructBLIP-7b model Dai et al. (2023a)). Experimental evaluation on established LVLM hallucination benchmarks shows that SSL outperforms existing decoding approaches, confirming its effectiveness and efficiency in hallucination reduction.
@@
  1. We leverage SAEs to identify semantic directions that are highly correlated with hallucinatory and faithful content in the representation space of LVLMs.
  2. We propose SSL, a plug-and-play method that injects factuality semantic directions during visual feature fusion to reinforce grounded content and suppresses hallucination directions during language generation to proactively mitigate hallucinatory outputs.
@@ 3 Method
In this work, we introduce SSL, a plug-and-play method for steering LVLMs. Our method consists of two principal components: semantic directions identification and steering LVLMs via SAE latents.
@@ 3.1 Semantic Directions Identification
Finally, to quantitatively evaluate the predictive power of SAE-derived directions in distinguishing hallucinatory from faithful samples, we design a set of classification experiments based on logistic regression. The model takes the latent activation values—either individually or in combination—as input features for a binary hallucination classification task. As illustrated in Figure 4, the results demonstrate that the latent semantic directions extracted by the SAE are discriminative (see Appendix B for more experimental details). Furthermore, combining the hall latent and faithful latent as input features yields further performance improvements. A more detailed discussion on clarifying the boundary between hallucinatory and faithful semantic directions can be found in Appendix C.
@@ 3.2 Steering LVLMs via SAE Latents
Input: Scaling factor $\gamma$; steering layer $l_{s}$; semantic directions $d_{\text{hall}}$, $d_{\text{faithful}}$; residual stream at layer $l_{s}$: $X_{l_{s},s},X_{l_{s},t},X_{l_{s},v},X_{l_{s},o}^{<t}]$
@@ 4.1 LVLMs
$\text{CHAIR}_{S}$ $\downarrow$ $\text{CHAIR}_{I}$ $\downarrow$ Avg.Len $\text{CHAIR}_{S}$ $\downarrow$ $\text{CHAIR}_{I}$ $\downarrow$ Avg.Len $\text{CHAIR}_{S}$ $\downarrow$ $\text{CHAIR}_{I}$ $\downarrow$ Avg.Len
@@ 4.2 Benchmarks
$\text{CHAIR}_{S}$ $=\frac{\left|\{\text{captions w/ hallucinatory objects}\}\right|}{\left|\{\text{total captions}\}\right|},$ (9)
$\text{CHAIR}_{I}$ $=\frac{\left|\{\text{hallucinatory objects}\}\right|}{\left|\{\text{total mentioned objects}\}\right|}.$
@@ 4.4 Implementation Details
We set $\gamma$ to $0.6$, $0.8$, and $0.1$ for LLaVA-NeXT, LLaVA-1.5, and InstructBLIP, respectively, to balance effective mitigation of hallucination with minimizing the invasiveness of state interventions. SSL is applied at $16^{\text{th}}$ layer for LLaVA-NeXT, $31^{\text{th}}$ layer for LLaVA-1.5 and $8^{\text{th}}$ layer for InstructBLIP. We faithfully replicate all baseline methods, implementing them based on their open-source codebases and configuring them according to the hyperparameters reported in the original papers. All experimental results are obtained under consistent base model, prompt, and generation parameter settings to ensure a fair comparison. For all methods involving beam search, we set the max_new_token to $512$ and the beam size to $5$.
@@ 5.1 Ablation Study
We conduct an ablation study on LLaVA-NeXT-8b to examine the impact of applying guidance at different layers. Figure 5 shows that the choice of guidance layer significantly affects model performance, setting $\gamma$ to $0.8$. For LLaVA-NeXT-8b, applying SSL at middle layers yields more effective mitigation of hallucinations, with layer $15$ achieving the best performance. For further analysis on steering layer selection and scaling factor $\gamma$ settings across different models, please refer to Appendix D and E.
@@ 5.2 Further Analysis
$\text{CHAIR}_{S}$ $\downarrow$ $\text{CHAIR}_{I}$ $\downarrow$
LLaVA-NeXT w/ ASP 26.36 6.32
@@
InstructBLIP w/ ASP 44.04 12.64
InstructBLIP w/ fixed $\alpha$ 45.76 13.12
@@
To further validate the effectiveness of the semantic directions we identified, we compare the CHAIR benchmark across three model architectures under three settings: original model state, SSL, and Reverse-SSL (see Appendix F for details on Reverse-SSL). As shown in Figure 6, across all three model architectures, applying Reverse-SSL significantly increases hallucinations, while standard SSL guidance consistently reduces hallucinations. To further clearly illustrate the effectiveness of SSL, we present several qualitative examples in Appendix G.
@@
$\text{CHAIR}_{S}$ $\downarrow$ $\text{CHAIR}_{I}$ $\downarrow$
LLaVA-1.5-7b 49.44 14.19
LLaVA-1.5-7b w/ SSL 40.88 12.30
LLaVA-Next-8b 29.60 8.03
LLaVA-Next-8b w/ SSL 26.36 6.32
Llama-3.2-11b-Vision-Instruct 31.40 7.40
Llama-3.2-11b-Vision-Instruct w/ SSL 28.92 6.69
Table 5: CHAIR results on MSCOCO dataset averaged over 5 random seeds. The best are indicated in bold.
Effect of Model Size on Method Performance.
To further analyze the impact of model size on our method, we conducted additional experiments on the 7B, 8B, and 11B models, evaluating SSL performance on both the CHAIR and POPE benchmarks, with results reported in Table 5 and Table 6. Specifically, for the 11B model, we use Llama-3.2-11b-Vision-Instruct released by Meta Llama, where SSL is applied with $\gamma=0.4$ at the $32^{\text{th}}$ layer. The results demonstrate that SSL effectively mitigates hallucinations across models of different sizes and consistently outperforms other baseline methods.
@@ 7 Conclusion
This work explores the relationship between the hallucination in LVLMs and their internal latent representations. We construct a residual stream dataset for hallucinatory and faithful object tokens, and use SAE to extract the semantic directions corresponding to hallucination and Factuality. Based on this insight, we propose SSL, a plug-and-play method that amplifies true semantics while suppressing potential hallucinations. Extensive experiments demonstrate that SSL outperforms existing methods. Furthermore, although the SAE was trained on LLaVA-Next, the semantic directions it extracted generalize well across different model architectures, further showcasing the potential of SAE in understanding and controlling the internal semantic representations of models.
@@
Currently, the only fully open-source multi-modal SAE is provided by LLM-Labs, trained on the $25^{\text{th}}$ layer of the LLaVA-Next 8b model. As a result, our study does not include a comparison of SAEs trained on other model architectures across different multimodal models. Future work can focus on training multi-modal SAEs on various architectures to investigate whether the findings from this study generalize across different models.
Acknowledgements
This work was supported by National Key R&D Program of China under Grant No.2024YFC3015501, also supported by the National Natural Science Foundation of China under Grant 62576094, and Grant U24A20322.
@@ Acknowledgements
$\text{Accuracy}\uparrow$ $\text{F1 score}\uparrow$
LLaVA-1.5-7b 83.86 84.99
LLaVA-1.5-7b w/ SSL 84.55 85.47
LLaVA-Next-8b 89.10 89.10
LLaVA-Next-8b w/ SSL 89.19 89.66
Llama-3.2-11b-Vision-Instruct 87.26 87.80
Llama-3.2-11b-Vision-Instruct w/ SSL 87.48 88.07
Table 6: POPE results averaged over popular, adversarial, and random splits. The best are indicated in bold.
@@ Appendix B A Set of Classification Experiments Based on Logistic Regression
Hall latent: The dimension hallhall with the highest correlation to hallucinated object words is selected, forming a one-dimensional feature:
@@
Faithful latent: The dimension faithfulfaithful most correlated with faithful object words is selected, forming:
@@
$\text{CHAIR}_{S}$ $\downarrow$ $\text{CHAIR}_{I}$ $\downarrow$ Avg.Len
@@
Appendix C Boundary Between Hallucinatory and Faithful Semantic Directions
First, we computed the activation frequency of each latent direction across hallucinatory and faithful samples. If a direction was significantly more active in hallucinatory samples, it was labeled as hallucinatory; conversely, it was labeled as faithful. Based on this criterion, we selected the top 128 directions with the largest activation differences between the two categories (excluding dead latents) for further analysis.
We then used a Support Vector Machine (SVM) classifier in the high-dimensional space to identify a separating decision boundary between the two categories, and applied Principal Component Analysis (PCA) to project the features into two dimensions for intuitive visualization.
Figure 10: Visualization of the linear SVM decision boundary in the 2D PCA-reduced latent space. Blue points represent faithful latent directions, while red points represent hallucinatory latent directions. The shaded regions indicate the SVM classification regions, and the black line shows the decision boundary.
As shown in Figure 10, although hallucinatory and faithful semantic directions appear as relatively tight clusters in the 2D projection space, the linear decision boundary learned by the SVM still effectively separates them. This apparent overlap is mainly due to PCA prioritizing the preservation of global variance rather than discriminative features, with much of the key class-distinguishing information compressed into higher-dimensional subspaces. Nevertheless, the presence of a clear linear boundary even after dimensionality reduction indicates that hallucinatory and faithful semantic directions are inherently separable, thereby validating the existence of a semantic boundary between the two groups.
Appendix D Choice of Steering Layer
@@
Figures 1112 and 13 present the results of ablation studies investigating the effect of introducing SSL at individual layers of the LLaVA-Next ($\gamma=0.6$), LLaVA-1.5 ($\gamma=0.8)$ and InstructBLIP ($\gamma=0.1$) models, respectively. For LLaVA-Next, we observe that applying SSL at the middle layers more effectively mitigates hallucinations, consistent with the results shown in Figure 5. For LLaVA-1.5, we observe that applying SSL at either the layer1 or deeper layers consistently mitigates hallucination. This observation aligns closely with findings reported by Zhang et al. (2024b), He et al. (2024) and Chen et al. (2025), who also found that layer1 or deeper layers interventions can significantly reduce hallucination in LLaVA-1.5. In contrast, for InstructBLIP, introducing SSL at shallow layers yields more substantial improvements, while deeper layer interventions contribute less noticeably to performance. We hypothesize that this is attributable to architectural and training differences in InstructBLIP, specifically, its shallow layers may already perform substantial cross-modal alignment early in the pipeline, making early-stage semantic guidance more impactful on overall generation quality. A deeper analysis of the layer-specific mechanisms in different multimodal architectures is left for future work.
Input: Scaling factor $\gamma$; steering layer $l_{s}$; semantic directions $d_{\text{hall}}$, $d_{\text{faithful}}$; residual stream at layer $l_{s}$: $X_{l_{s},s},X_{l_{s},t},X_{l_{s},v},X_{l_{s},o}^{<t}]$
@@
Appendix E Choice of Scaling Factor
Tables 78, and 9 report the ablation results on the effect of the scaling factor $\gamma$ in the SSL. For the LLaVA series of models, setting $\gamma$ to 0.6 or 0.8 effectively reduces hallucinations, indicating that moderate levels of semantic intervention are beneficial. However, when $\gamma>1.0$, the supervision becomes overly aggressive, disrupting the model’s behavior and leading to abnormal hallucination metrics. In contrast, for InstructBLIP, larger values similarly result in performance degradation. These findings highlight the importance of carefully calibrating the intensity of semantic guidance to balance model control and generation quality across different architectures.
$\text{CHAIR}_{S}$ $\downarrow$ $\text{CHAIR}_{I}$ $\downarrow$ Avg.Len